# Bending-strain effects in conventional superconductors and superconducting junctions

### v1

@@ Appendix A Pair amplitudes
$\dot{\mathcal{P}}_{i,0}^{s}(\tau=0)$ $=$ $\frac{-i}{N_{b}}\sum_{\pm,n,k}^{{}^{\prime}}\pm\mathcal{E}_{n,k}\bigg{[}\big{(}x^{*}_{i\pm 1,n,k}u_{i,n,k}-w^{*}_{i\pm 1,n,k}v_{i,n,k}\big{)}f(-2\mathcal{E}_{n,k})$ (37)
@@
$\mathcal{P}^{b}_{i,0}(\tau)$ $=$ $\frac{-i}{N_{b}}\sum_{n,k}^{{}^{\prime}}\sin(ka)\left(x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\right)\left(f(2\mathcal{E}_{n,k})e^{2i\mathcal{E}_{n,k}\tau}+f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}\right),$ (38)
$\dot{\mathcal{P}}_{i,0}^{b}(\tau=0)$ $=$ $\frac{-1}{N_{b}}\sum_{n,k}^{{}^{\prime}}\mathcal{E}_{n,k}\sin(ka)\left(x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\right)\tanh(\beta\mathcal{E}_{n,k}),$ (39)
$\mathcal{P}_{i,\uparrow\downarrow}^{s}$ $=$ $\frac{1}{2N_{b}}\sum_{n,k,\pm}^{{}^{\prime}}\pm\biggl{[}\left(x^{*}_{i\pm 1,n,k}u_{i,n,k}+w^{*}_{i\pm 1,n,k}v_{i,n,k}\right)f(-2\mathcal{E}_{n,k})$ (40)
@@
$\mathcal{P}_{i,\uparrow\uparrow}^{b}$ $=$ $\frac{-2i}{N_{b}}\sum_{n,k}^{{}^{\prime}}w^{*}_{i,n,k}u_{i,n,k}\sin(ka)\tanh(\beta\mathcal{E}_{n,k}),$ (41)
$\mathcal{P}_{i,\downarrow\downarrow}^{b}$ $=$ $\frac{-2i}{N_{b}}\sum_{n,k}^{{}^{\prime}}x^{*}_{i,n,k}v_{i,n,k}\sin(ka)\tanh(\beta\mathcal{E}_{n,k}).$ (42)
Appendix B Odd-frequency amplitudes
@@
The odd-frequency order parameters are influenced by the relative time coordinate $\tau$, which is scaled by $\hbar/t$ and may appear arbitrary. We have found that the amplitudes remain within the same order of magnitude for relative times that are within the same order. They are also comparable to the derivative evaluated at zero. In Fig. B.1, we have plotted the order parameters at $\tau=1$, along with their derivatives evaluated at $\tau=0$.
@@
Figure B.1: The odd-frequency order parameters that are finite at the superconductor-normal metal interface along with their relative time derivatives. The first column is evaluated at $\tau=1$, and their derivatives in the second column at $\tau=0$.
Appendix C Spin-current $0-\pi$ oscillations with length
@@
In an SNS junction, the binormal component of the spin-current in the tangential direction, $J^{b}$, undergoes $0-\pi$ oscillations as a function of both the applied strain to the superconductors, as seen in Fig. 5, and as a function of the length of the normal metal $N_{N}$. These $0-\pi$-transitions as a function of length are shown in Fig. C.2, where the transition in strain can also be seen for $N_{N}=3,6$. Further changes in sign occur at other lengths for higher strain (see Fig. 5).
@@
Figure C.2: The binormal component of the spin-current exhibits $0-\pi$ oscillations as a function of increasing number of lattice sites in the normal metal, $N_{N}$. The values are taken from the center of the normal metal. We have used $N_{S}=40$, $N_{b}=99$, and $N_{n}=3$ sites.
Appendix D Comparison of curvatures, strain and SOC-energy
@@
We take the strain-induced SOC Salamone et al. (2022) $\alpha_{N}=a_{N}\kappa$ to be of the order $\alpha_{N}\approx\kappa$. For specificity, we consider a BCC lattice Niobium film with $a=1.65$ González-Pedreros et al. (2021), $m^{*}=2.14m_{e}$, $|k_{F}|=1.18$-1,Prozorov et al. (2022) and the gap $|\Delta|=2.32$ meV Bonnet et al. (1967). In Table C.1, we then compare the SOC energy $\Delta_{\text{soc}}=\hbar^{2}\kappa|k_{F}|/m^{*}$ with the gap for different values of strain, and show how this relates to the geometric curvature $\kappa=1/R$, where $R$ is the radius of curvature, for three different film thicknesses, $N_{n}=(3,6,9)$. Note that since the SOC strength scales with the curvature, and the strain scales with the curvature and thickness, the effective SOC strength scales with strain and the inverse of the thickness. This is apparent in the Fig. 3, where we see a shift along strain-axis for increasing $N_{n}$.
@@
Table C.1: In the table above are different thicknesses of Niobium ($N_{n}$ layers). We give the curvature $\kappa$ and the resulting strain $\epsilon_{ss}$ (%). In the last column we compare the SOC energy $\Delta_{\text{soc}}$ with the superconducting gap $\Delta$.

### v2

@@ V Pair amplitudes
$\dot{\mathcal{P}}_{i,0}^{s}(\tau=0)$ $=$ $\frac{-i}{N_{b}}\sum_{\pm,n,k}^{{}^{\prime}}\pm\mathcal{E}_{n,k}\bigg{[}\big{(}x^{*}_{i\pm 1,n,k}u_{i,n,k}-w^{*}_{i\pm 1,n,k}v_{i,n,k}\big{)}f(-2\mathcal{E}_{n,k})$ (S14)
@@
$\mathcal{P}^{b}_{i,0}(\tau)$ $=$ $\frac{-i}{N_{b}}\sum_{n,k}^{{}^{\prime}}\sin(ka)\left(x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\right)\left(f(2\mathcal{E}_{n,k})e^{2i\mathcal{E}_{n,k}\tau}+f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}\right),$ (S15)
$\dot{\mathcal{P}}_{i,0}^{b}(\tau=0)$ $=$ $\frac{-1}{N_{b}}\sum_{n,k}^{{}^{\prime}}\mathcal{E}_{n,k}\sin(ka)\left(x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\right)\tanh(\beta\mathcal{E}_{n,k}),$ (S16)
$\mathcal{P}_{i,\uparrow\downarrow}^{s}$ $=$ $\frac{1}{2N_{b}}\sum_{n,k,\pm}^{{}^{\prime}}\pm\biggl{[}\left(x^{*}_{i\pm 1,n,k}u_{i,n,k}+w^{*}_{i\pm 1,n,k}v_{i,n,k}\right)f(-2\mathcal{E}_{n,k})$ (S17)
@@
$\mathcal{P}_{i,\uparrow\uparrow}^{b}$ $=$ $\frac{-2i}{N_{b}}\sum_{n,k}^{{}^{\prime}}w^{*}_{i,n,k}u_{i,n,k}\sin(ka)\tanh(\beta\mathcal{E}_{n,k}),$ (S18)
$\mathcal{P}_{i,\downarrow\downarrow}^{b}$ $=$ $\frac{-2i}{N_{b}}\sum_{n,k}^{{}^{\prime}}x^{*}_{i,n,k}v_{i,n,k}\sin(ka)\tanh(\beta\mathcal{E}_{n,k}).$ (S19)
We have used that the time-dependent operators are given by $c_{\bm{i},\sigma}(\tau)=e^{i\mathcal{H}\tau}c_{\bm{i},\sigma}e^{-i\mathcal{H}\tau}$, where $\mathcal{H}$ is the Hamiltonian and $\tau$ the relative time coordinate. The expectation value of the number operator in Bogoliubov quasiparticle operators becomes ${\big{<}{\gamma^{\dagger}_{n,k}\gamma_{n^{\prime},k}}\big{>}=f(2\mathcal{E}_{n,k})\delta_{n,n^{\prime}}}$, where $f(2\mathcal{E}_{n,k})$ is the Fermi-Dirac distribution, and $\beta$ is the inverse temperature.
VI Odd-frequency amplitudes
@@
The odd-frequency order parameters are influenced by the relative time coordinate $\tau$, which is scaled by $\hbar/t$ and may appear arbitrary. We have found that the amplitudes remain within the same order of magnitude for relative times that are within the same order. They are also comparable to the derivative evaluated at zero. In Fig. S1, we have plotted the order parameters at $\tau=1$, along with their derivatives evaluated at $\tau=0$.
@@
Figure S1: The odd-frequency order parameters that are finite at the superconductor-normal metal interface along with their relative time derivatives. The first column is evaluated at $\tau=1$, and their derivatives in the second column at $\tau=0$.
VII Spin-current $0-\pi$ oscillations with length
@@
In an SNS junction, the binormal component of the spin-current in the tangential direction, $J^{b}$, undergoes $0-\pi$ oscillations as a function of both the applied strain to the superconductors, as seen in Fig. (5) in the main text, and as a function of the length of the normal metal $N_{N}$. These $0-\pi$-transitions as a function of length are shown in Fig. S2, where the transition in strain can also be seen for $N_{N}=3,6$. Further changes in sign occur at other lengths for higher strain (see Fig. (5) in the main text).
@@
Figure S2: The binormal component of the spin-current exhibits $0-\pi$ oscillations as a function of increasing number of lattice sites in the normal metal, $N_{N}$. The values are taken from the center of the normal metal. We have used $N_{S}=40$, $N_{b}=99$, and $N_{n}=3$ sites.
VIII Comparison of curvatures, strain and SOC-energy
@@
We take the strain-induced SOC Salamone et al. (2022) $\alpha_{N}=a_{N}\kappa$ to be of the order $\alpha_{N}\approx\kappa$. For specificity, we consider a BCC lattice Niobium film with $a=1.65$González-Pedreros et al. (2021), $m^{*}=2.14m_{e}$, $|k_{F}|=1.18$-1,Prozorov et al. (2022) and the gap $|\Delta|=2.32$ meV Bonnet et al. (1967). In Table 1, we then compare the SOC energy $\Delta_{\text{soc}}=\hbar^{2}\kappa|k_{F}|/m^{*}$ with the gap for different values of strain, and show how this relates to the geometric curvature $\kappa=1/R$, where $R$ is the radius of curvature, for three different film thicknesses, $N_{n}=(3,6,9)$. Note that since the SOC strength scales with the curvature, and the strain scales with the curvature and thickness, the effective SOC strength scales with strain and the inverse of the thickness. This is apparent in Fig. (3) in the main text, where we see a shift along strain-axis for increasing $N_{n}$.
@@
Table 1: In the table above are different thicknesses of Niobium ($N_{n}$ layers). We give the curvature $\kappa$ and the resulting strain $\epsilon_{ss}$ (%). In the last column we compare the SOC energy $\Delta_{\text{soc}}$ with the superconducting gap $\Delta$.